OpenStudy (anonymous):

(x^2-6x)^2-35(x^2-6x)-200=0

myininaya (myininaya):

solve u^2-35u-200=0 find two factors of -200 that have product -200 and have sum -35 -200=-40(5) -35=-40+5 now replace bu=-35u with -40u+5u like u^2-40u+5u-200=0 now factor by grouping

myininaya (myininaya):

let me see the step after this one

myininaya (myininaya):

show me whenever you like